OpenStudy (anonymous):

how to factor x^3 - y^3 show steps please

OpenStudy (anonymous):



OpenStudy (anonymous):

I always thought it was an identity and no steps ....just an answer\[x^3-y^3=(x-y)(x^2+xy+y^2)\]

OpenStudy (anonymous):

the identity has to come from somewhere

OpenStudy (anonymous):

Choose your god !

OpenStudy (anonymous):

You could possibly use algebraic long division with the first factor known to get the second factor.

OpenStudy (anonymous):

choose your god? and im not sure how to go about doing long division on this

OpenStudy (anonymous):

we can open \[(x+y)^3=x^3+3xy^2+3x^2y+y^3\] \[(x+y)^3-3xy(x+y)=x^3+y^3\] \[(x+y)((x+y)^2-3xy)=x^3+y^3\] \[(x+y)(x^2+2xy-3xy+y^2)=x^3+y^3\] \[(x+y)(x^2-xy+y^2)=x^3+y^3\]

OpenStudy (anonymous):

this is for plus but you can do for -

OpenStudy (anonymous):

Very nice work Jonask

OpenStudy (anonymous):

alright thanks a lot

OpenStudy (anonymous):

can you do -

OpenStudy (anonymous):

x^3-y^3

OpenStudy (anonymous):

(x-y)^3